OpenStudy (anonymous):

Three-fifths of the students in the school orchestra play the violin. if 2/3 of the violin players are girls and 24 girls play violin, how many students are in the orchestra? BTW im study for the acts this is one thing i need to know how to do

OpenStudy (anonymous):

You should start by figuring out how many violin players there are.

OpenStudy (anonymous):

how do i go about doing that, division?

OpenStudy (anonymous):

2/3 = 24 girls

OpenStudy (anonymous):

so i would have to find 2/3 of 3/4 right?

OpenStudy (whpalmer4):

My ten year old just figured it out and is in 5th grade. Here's what he did: 24 girls play the violin. 24 is 2/3 of what number?

OpenStudy (whpalmer4):

if 24 is 2/3, then half of 24 is 1/3. half of 24 is 12. 3*12 is 36. so 36 kids play the violin. 36 is 3/5 of all of the kids in the orchestra. figure out how many kids are in 1/5 if 36 is 3/5. add two more fifths to 36 to get the total in the orchestra

OpenStudy (whpalmer4):

working backwards like he did is often a very effective way to solve these problems.

OpenStudy (anonymous):

^ couldn't have said it better myself :P

OpenStudy (anonymous):

it said it was 60

OpenStudy (anonymous):

You're right

OpenStudy (whpalmer4):

Yes. Let's check (always a good plan!): 60 kids, 3/5 play violin, so 3/5*60 = 180/5 = 36 kids play violin. 2/3 of violinists are girls: 2/3*36 = 72/3 = 24 girls playing violin. Everything checks out.